Grade 12th passPhysical Chemistry

  1. the potential energy of smple harmonic oscillater of mass 1kg in its mean position is 3j.and its total energy is 7j. it its time period is 2root 2×10^-2 find its amplitude

kalyan
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
K.E at the centre= 7-3 = 4J
According to the question :
 1/2 K A^2 = 4j
k = kinetic energy and A = amplitude
 k = 8/ A^2 
T = 2*pie sqrt(m/k)
2*pie*sqrt(1/50) 
k = 50 
A^2 = 8/50 
A= 0.4
Hope this helps 
Good luck